For project A, the expected investment is $ 1 M and annual Cash Flows are 300K. For project B, the investment is $ 2 M and cash flows are 500K. Economic life for each project is 10 years. Projects are mutually exclusive. a. What is the incremental discounted rate of return? 1. 30% ii. 20% iii. 15% iv, 25%
If the minimum attractive interest rte is 10% for the above projects, at what year project B will be more attractive. i.5 iv. 9 v. never
Answer #1

Project A

Project B

Incremental CF of B – A

Initial Investment

-1,000,000

-2,000,000

-1,000,000

Cash Flows

300,000

500,000

200,000

a. Incremental discounted rate of return

Using the trial and error method

Life of the projects = 10 years

Let the rate of interest is 14%

At 14% the NPW is

NPW at 14% = -1,000,000 + 200,000 (P/A, 14%, 10)

NPW at 14% = -1,000,000 + 200,000 (5.2161) = 43,220

Since the PW is positive increase the rate of interest to 16% for getting a negative NPW

NPW at 16%

NPW at 16% = -1,000,000 + 200,000 (P/A, 16%, 10)

NPW at 16% = -1,000,000 + 200,000 (4.8332) = -33,360

Using interpolation

Rate of return = 14% + [43,220 – 0 ÷ 43,220 – (-33,360)]*2% = 15.1%

Incremental Discount rate of return will be 15%

Answer – iii. 15%

b. If MARR is 10%, at what year project B will be attractive.

Calculating the discounted payback period of Project B

Year

CF

PV Factor

DCF

CCF

0

($2,000,000)

1

($2,000,000)

($2,000,000)

1

$500,000

0.91

$454,545.45

($1,545,454.55)

2

$500,000

0.83

$413,223.14

($1,132,231.40)

3

$500,000

0.75

$375,657.40

($756,574)

4

$500,000

0.68

$341,506.73

($415,067.28)

5

$500,000

0.62

$310,460.66

($104,606.62)

6

$500,000

0.56

$282,236.97

$177,630.35

Discounted Payback Period = 5 + [-104,606.62 – 0 ÷ -104,606.62 – (177,630.35)]*1 = 5.37 years

Answer – i. 5 years